# Developmentally-plausible Working Memory Shapes a Critical Period for Language Acquisition

### v2

@@ Abstract
[…]
@@ 1 Introduction
An important theoretical framework for understanding the efficiency of human language acquisition is the Critical Period Hypothesis (CPH) Lenneberg (1967). The CPH posits that there is a specific period during which language can be acquired efficiently, and that this ability diminishes thereafter. Various studies, including cases of limited first language ($\mathrm{L}_{1}$) exposure during childhood and age-related effects on second language ($\mathrm{L}_{2}$) acquisition, support the existence of a critical period (CP) Fromkin et al. (1974); Curtiss (1977); Johnson and Newport (1989). However, the reasons why children acquire language more efficiently than adults remain partially unresolved. One compelling explanation for the CP in $\mathrm{L}_{1}$ acquisition is the Less-is-More Hypothesis Newport (1990), which argues that children’s cognitive limitations (e.g., working memory capacity and attentional scope) are advantageous for language learning. According to this hypothesis, children’s limited processing capacities enable them to efficiently extract fundamental patterns and structures (e.g., grammatical rules) from linguistic input, whereas adults, with their greater cognitive capacities, are more likely to be distracted by complex information, thereby hindering rule acquisition.
@@
Inspired by the “Less-is-More” hypothesis, we use language models (LMs) to study the CP for language acquisition, focusing on $\mathrm{L}_{1}$ acquisition and investigating whether integrating human cognitive developmental characteristics, particularly the developmental properties of working memory (Figure 1), into LMs can facilitate efficient language acquisition. Specifically, we propose a method for incorporating the exponential increase in working memory capacity that corresponds to the CP into LMs and analyze its impact on learning efficiency. Using a GPT-2 model Radford et al. (2019) trained on a Child-Directed Speech (CDS) dataset Huebner and Willits (2021), we conduct evaluation experiments with Zorro Huebner et al. (2021), a targeted syntactic evaluation benchmark specialized for CDS. The results demonstrate that a cognitively plausible model, which initially restricts working memory and gradually relaxes this constraint exponentially as training progresses, outperforms models without memory constraints or with static memory constraints. These findings provide new insights into designing data-efficient LMs, contributing to the field of natural language processing, while also offering indirect evidence supporting the role of the developmental characteristics of working memory as the underlying mechanism of the CPH in human language acquisition, contributing to the field of cognitive science.
@@ 2.2 The Role of Language Models in Acquisition Theories
In recent years, computational models have played a crucial role in elucidating the mechanisms of language acquisition. These models enable controlled investigations of learning mechanisms and environments, which are difficult to achieve with human participants, and they are used to test theoretical claims such as the “poverty of the stimulus” Clark and Lappin (2011). For instance, McCoy et al. (2020), Wilcox et al. (2024), and Warstadt et al. (2023) have employed LMs to directly test hypotheses about language acquisition, demonstrating that such models can provide proof-of-concept evidence for learnability. These studies have attracted attention as efforts to deepen theoretical discussions on language acquisition through computational modeling, including research on the CP.
Constantinescu et al. (2025) investigated CP phenomena in $\mathrm{L}_{2}$ acquisition and $\mathrm{L}_{1}$ attrition, [^1] The phenomenon in which earlier cessation of $\mathrm{L}_{1}$ exposure increases the likelihood of $\mathrm{L}_{1}$ forgetting. assuming a shared underlying mechanism for CP effects across $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$. They simulated $\mathrm{L}_{2}$ exposure at varying ages to examine how LMs differ from human learners, finding that LMs do not naturally exhibit CP effects. To artificially induce such effects, they employed Elastic Weight Consolidation Kirkpatrick et al. (2017), a regularization method for mitigating catastrophic forgetting, thereby mimicking a maturational decline in plasticity. Their findings suggest that CP effects are not an inevitable outcome of statistical learning but may instead involve innate mechanisms.
@@
While this study shares the broader objective of enhancing the cognitive plausibility of LMs as models of human language acquisition, it differs from Constantinescu et al. (2025) in both focus and methodology. Rather than modeling CP effects through dataset manipulation or post-CP plasticity constraints, this study explicitly addresses the developmental processes unfolding during the CP itself. Specifically, we integrate a mechanism to simulate the progressive growth of working memory capacity throughout the CP, a factor considered crucial for $\mathrm{L}_{1}$ acquisition but previously unmodeled in LM-based research. By incorporating developmental constraints, this study aims to provide a more fine-grained computational model of early $\mathrm{L}_{1}$ acquisition and its cognitive underpinnings, advancing the developmental plausibility of LMs.
3 Language Model with Developmentally-plausible Working Memory
@@ 3.1 Modeling Developmental Trajectory of Human Working Memory
This modeling approach is justified for several reasons. First, the horizontal asymptote inherent in the exponential function accurately represents the biological ceiling of adult working memory capacity. Second, the rapid initial increase observed during early childhood is consistent with the steep growth predicted by this exponential form. Finally, alternative models, such as logarithmic or linear growth, fail to account for both the early rapid development and the eventual plateau: logarithmic models imply unbounded growth, while linear models oversimplify the deceleration phase. Thus, the exponential model $y=b-a^{x}$ offers a concise and biologically plausible representation of the developmental trajectory of human working memory, aligning well with observed patterns and theoretical considerations.
@@ 3.2 Integrating Human Working Memory into Language Models
This definition establishes a direct relationship between the dynamically decaying slope $m_{t}$ and the model’s working memory capacity $w_{t}$. As $m_{t}$ decreases exponentially over time, $w_{t}$, representing working memory, grows correspondingly, allowing the model to retain broader contextual information as training progresses. By mimicking this developmentally plausible growth of working memory, the model prioritizes attention to short-range dependencies during the early stages of training, gradually shifting its focus to long-range dependencies as training progresses.
Furthermore, a key distinction between ALiBi and DynamicLimit-Exp lies in how the slope $m$ is assigned across attention heads. While ALiBi applies a fixed per-head bias, enforcing a predetermined recency bias throughout training, DynamicLimit-Exp instead shares the slope $m$ across all heads. This ensures that the model maintains a globally coherent bias that evolves dynamically over the course of training. In other words, ALiBi imposes a head-specific static recency bias, whereas DynamicLimit-Exp introduces a dynamically changing proximity bias that governs the entire learning schedule. This shift enables the model to more accurately simulate the adaptive nature of human working memory development, potentially capturing the CP of cognitive maturation.
@@ 4.1 Configurations
Table 1: Accuracy (%) of models trained on AO-CHILDES dataset. Overall represents the macro average of the scores across all grammar items.
@@
Table 2: Accuracy (%) of models trained on Wikipedia dataset. Overall represents the macro average of the scores across all grammatical items.
@@
Furthermore, to determine whether the CP effect stems from exposure to specific linguistic stimuli, such as CDS, or from the model’s cognitive developmental properties independent of input, we conducted a complementary experiment using Wikipedia (written language, adult-oriented) as training data. Following Huebner et al. (2021), 500,000 sentences were randomly sampled from the English Wikipedia corpus. We used the latest version of Wikipedia, as of January 2025,  [^3] https://dumps.wikimedia.org/enwiki/latest/enwiki-latest-pages-articles.xml.bz2 and preprocessed it using WikiExtractor. [^4] https://github.com/attardi/wikiextractor
@@
We evaluate the grammatical abilities of these models using a developmentally inspired targeted syntactic evaluation benchmark, Zorro Huebner et al. (2021). Zorro is designed for assessing the syntactic and grammatical knowledge of LMs in child-directed language and consists of 13 mid-level categories and 23 subcategories. Each subcategory contains 2,000 sentence pairs, with one grammatically acceptable and one unacceptable sentence per pair. Below is an example of a minimal pair from the “Subject-verb agreement (S-v agr)” category: [^5] See Appendix B for the full list of grammatical categories.
@@
By inputting both the acceptable and unacceptable sentence into the model and calculating the proportion of pairs where the model assigns a higher probability to the acceptable sentence, we obtain the grammaticality judgment score (Accuracy). In this study, we report scores for each mid-level category (henceforth, grammatical items) as well as their macro-average.
@@
Table 3: Performance difference when changing the direction of the cognitive constraints in DynamicLimit-Exp
@@ 4.3 Results
Table 2 presents the accuracy of models trained on Wikipedia, showing trends similar to those observed in Table 1, where the models were trained on AO-CHILDES. Specifically, DynamicLimit-Linear and DynamicLimit-Exp outperform NoLimit and StaticLimit in overall accuracy, with DynamicLimit-Exp achieving the highest performance, further supporting the efficacy of incorporating developmental working memory constraints. These findings suggest that the CP effect does not depend solely on exposure to specific linguistic stimuli (e.g., CDS) but rather on the learning algorithm itself, which mirrors human cognitive development.
This result aligns with existing research Feng et al. (2024), which has reported that child language input is not uniquely valuable for training LMs. This finding suggests that our method is applicable to LLM pretraining, as they typically use non-CDS datasets such as Common Crawl and Wikipedia Touvron et al. (2023a).
(a) NoLimit
(b) DynamicLimit-Exp
@@ 5.1 Testing the “Less-is-more” Hypothesis with Reversed Cognitive Constraints
[…]
Mark fixed one worn canal, and Roger
  fixed more.
*Mark fixed one canal, and Roger fixed
5.2 Development of Feature Extraction Capabilities
Figure 3 visualizes the clustering structure of final-layer embeddings using t-SNE van der Maaten and Hinton (2008) for FILLER.GAP, a grammatical items where gradual memory expansion yielded significant performance improvements in both AO-CHILDES and Wikipedia datasets, as highlighted in the previous results (§4.3 and §5.1). In NoLimit (Figure 3a), the embedding clusters initially expand between Epoch 1 and Epoch 5, but by Epoch 10, they appear to contract and overlap more, suggesting a stagnation in representation learning. The clusters become less distinguishable, which may indicate a loss of diversity in the learned representations. In contrast, DynamicLimit-Exp (Figure 3b) maintains a more structured and progressive evolution of embeddings. The clusters remain well-separated throughout training, with clear distinctions between different epochs. This suggests that the model continuously refines its representations without excessive compression, preserving the diversity necessary for robust generalization.
To quantitatively analyze these differences, Table 4 reports key statistical measures, including entropy (distribution diversity) and mean Euclidean distance (inter-cluster separation). [^8] The appendix C shows how to calculate each measure. Regarding entropy, NoLimit shows a decreasing trend, reflecting reduced distribution diversity and potential over-clustering. In contrast, DynamicLimit-Exp preserves consistently higher entropy, indicating a balanced representation that avoids excessive compression. For mean Euclidean distance, NoLimit undergoes substantial change between Epoch 1 and Epoch 5 but stagnates thereafter, suggesting limited refinement. DynamicLimit-Exp, however, maintains large distances across epochs, reflecting continuous structural reorganization.
The differences also highlight the role of isotropy. NoLimit exhibits increasing anisotropy, with embedding clusters becoming overly compact by Epoch 10, which may hinder generalization. In contrast, DynamicLimit-Exp maintains a more isotropic distribution, as indicated by stable entropy, allowing for more flexible and structured representation learning. These findings align with recent work on syntactic smoothing, which suggests that reducing anisotropy enhances the ability to generalize across linguistic contexts Diehl Martinez et al. (2024). Thus, the increased isotropy observed in DynamicLimit-Exp provides strong evidence that gradual memory expansion facilitates structured representation learning and syntactic generalization. [^9] We also analyzed CASE, which exhibited the same trend as FILLER.GAP (as shown in Appendix D).

### v3

@@ Abstract
Masato Mita  and Ryo Yoshida  and Yohei Oseki The University of Tokyo {mita, yoshiryo0617, oseki}@g.ecc.u-tokyo.ac.jp
@@ 1 Introduction
An important theoretical framework for understanding the efficiency of human language acquisition is the Critical Period Hypothesis (CPH) Lenneberg (1967). The CPH posits that there is a specific period during which language can be acquired efficiently and that this ability diminishes thereafter. Various studies, including cases of limited exposure to first language ($\mathrm{L}_{1}$) during childhood and age-related effects on second language ($\mathrm{L}_{2}$) acquisition, support the existence of a critical period (CP) Fromkin et al. (1974); Curtiss (1977); Johnson and Newport (1989). However, the reasons why children acquire language more efficiently than adults remain partially unresolved. A compelling explanation is the Less-is-More Hypothesis Newport (1990), which attributes the superior learning of children to limited cognitive resources such as working memory. According to this hypothesis, children’s limited processing capacities enable them to efficiently extract fundamental patterns and structures (e.g., grammatical rules) from linguistic input, whereas adults, with their greater cognitive capacities, are more likely to be distracted by complex information, thereby hindering rule acquisition.
This hypothesis not only offers a compelling account of human learning but also has implications for how we design artificial systems. Language is not an arbitrary object of learning but a cultural artifact shaped under cognitive constraints. A growing body of work suggests that its structural properties reflect pressures for efficient communication—that is, to maximize informational content while minimizing cognitive effort in production and comprehension under human limitations (Zipf, 1949; Jaeger and Tily, 2011; Christiansen and Chater, 2016; Kemp et al., 2018; Fedorenko et al., 2024). Over generations, language has likely evolved to be learnable by agents with limited memory and processing capacity. From this perspective, incorporating such constraints into language models (LMs) is not merely an act of mimicking human limitations, but a theoretically grounded way to introduce an inductive bias that aligns with the nature of the target: language shaped by cognitively bounded agents. Learning under such constraints may help LMs acquire representations better suited to natural language. [^1] Futrell and Mahowald (2025) for an alternative view suggesting that, given the empirical success of machine learning, effective learning may not require cognitively inspired constraints or inductive biases.
Inspired by the Less-is-More hypothesis, we use LMs to study the CP for language acquisition, focusing on $\mathrm{L}_{1}$ acquisition and investigating whether integrating human cognitive developmental characteristics, particularly the developmental properties of working memory (Figure 1), into LMs can facilitate efficient language acquisition. Specifically, we propose a method for incorporating the exponential increase in working memory capacity that corresponds to the CP into LMs and analyze its impact on learning efficiency. Using a GPT-2 model Radford et al. (2019) trained on a Child-Directed Speech (CDS) dataset Huebner and Willits (2021), we conduct evaluation experiments with Zorro Huebner et al. (2021), a targeted syntactic evaluation benchmark specialized for CDS. The results demonstrate that a cognitively plausible model, which initially restricts working memory and gradually relaxes this constraint exponentially as training progresses, outperforms models without memory constraints or with static memory constraints. These findings provide new insights into designing data-efficient LMs, contributing to the field of natural language processing, while also offering indirect evidence supporting the role of the developmental characteristics of working memory as the underlying mechanism of the CPH in human language acquisition, contributing to the field of cognitive science.
@@ 2.2 The Role of Language Models in Acquisition Theories
In recent years, computational models have played a crucial role in elucidating the mechanisms of language acquisition. These models enable controlled investigations of learning mechanisms and environments, which are difficult to achieve with human participants, and they are used to test theoretical claims such as the “poverty of the stimulus” Clark and Lappin (2011). For instance, McCoy et al. (2020), Wilcox et al. (2024), and Warstadt et al. (2023) have employed LMs to directly test hypotheses about language acquisition, demonstrating that such models can provide proof-of-concept evidence for learnability. These studies have attracted attention as efforts to deepen theoretical discussions on language acquisition through computational modeling.
While Transformer-based LMs are not cognitive models in a strict sense, they are widely adopted in acquisition research as abstract “model learners” Warstadt and Bowman (2022). Rather than replicating the full complexity of human cognition, these models are used to investigate the role of specific biases by selectively adding or removing them. This approach allows researchers to assess whether certain linguistic phenomena can be acquired purely through statistical learning or require inductive constraints, thereby testing the necessity of such biases. When a model fails to acquire a phenomenon in the absence of a particular bias, but succeeds once the bias is introduced, it offers at least weak evidence for the bias’s relevance in human language acquisition McCoy et al. (2020). Our study follows this reverse-engineering paradigm Dupoux (2018), using LMs not as literal simulations of human learners, but as controlled testbeds for cognitive hypotheses.
@@
Constantinescu et al. (2025) investigated CP phenomena in $\mathrm{L}_{2}$ acquisition and $\mathrm{L}_{1}$ attrition, [^2] The phenomenon in which earlier cessation of $\mathrm{L}_{1}$ exposure increases the likelihood of $\mathrm{L}_{1}$ forgetting. assuming a shared underlying mechanism for CP effects across $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$. They simulated $\mathrm{L}_{2}$ exposure at varying ages to examine how LMs differ from human learners, finding that LMs do not naturally exhibit CP effects. To artificially induce such effects, they employed Elastic Weight Consolidation Kirkpatrick et al. (2017), a regularization method for mitigating catastrophic forgetting, thereby mimicking a maturational decline in plasticity. Their findings suggest that CP effects are not an inevitable outcome of statistical learning but may instead involve innate mechanisms. While this study shares the broader objective of enhancing the cognitive plausibility of LMs as models of human language acquisition, it differs from Constantinescu et al. (2025) in both focus and methodology. Rather than modeling CP effects through dataset manipulation or post-CP plasticity constraints, this study explicitly addresses the developmental processes unfolding during the CP itself. Specifically, we integrate a mechanism to simulate the progressive growth of working memory capacity throughout the CP, a factor considered crucial for $\mathrm{L}_{1}$ acquisition but previously unmodeled in LM-based research. By incorporating developmental constraints, this study aims to provide a more fine-grained computational model of early $\mathrm{L}_{1}$ acquisition and its cognitive underpinnings, advancing the developmental plausibility of LMs.
3 Critical Period-inspired Language Model
@@ 3.1 Modeling Developmental Trajectory of Human Working Memory
This modeling approach is justified for several reasons. First, the horizontal asymptote inherent in the exponential function accurately represents the biological ceiling of adult working memory capacity. Second, the rapid initial increase observed during early childhood is consistent with the steep growth predicted by this exponential form. Finally, alternative models, such as logarithmic or linear growth, fail to account for both the early rapid development and the eventual plateau: logarithmic models imply unbounded growth, while linear models oversimplify the deceleration phase. Thus, the exponential model $y=b-a^{x}$ offers a concise and biologically plausible representation of the developmental trajectory of human working memory, aligning well with observed patterns and theoretical considerations. [^3] ACT-R Anderson and Milson (1989) suggests that working memory decays exponentially in language processing, while we propose that working memory grows exponentially in language acquisition, but whether the shared exponential function between language processing and acquisition is a coincidence remains to be investigated in future.
@@ 3.2 Integrating Human Working Memory into Language Models
This definition links the dynamically decaying slope $m_{t}$ to the model’s working memory capacity $w_{t}$: as $m_{t}$ decreases exponentially, $w_{t}$ grows, enabling broader contextual retention over time. By simulating this developmental trajectory, the model initially focuses on short-range dependencies and gradually attends to longer ones.
@@ 4.1 Configurations
Table 1: Accuracy (%) of models trained on AO-CHILDES dataset. and indicate items where DynamicLimit-Exp performed significantly better or worse than NoLimit, respectively (z-test for proportions, p < 0.05).
@@
Table 2: Accuracy (%) of models trained on Wikipedia dataset. and indicate items where DynamicLimit-Exp performed significantly better or worse than NoLimit, respectively (z-test for proportions, p < 0.05).
@@
Furthermore, to determine whether the CP effect stems from exposure to specific linguistic stimuli, such as CDS, or from the model’s cognitive developmental properties independent of input, we conducted a complementary experiment using Wikipedia (written language, adult-oriented) as training data. Following Huebner et al. (2021), 500,000 sentences were randomly sampled from the English Wikipedia corpus. We used the latest version of Wikipedia, as of January 2025, [^5] https://dumps.wikimedia.org/enwiki/latest/enwiki-latest-pages-articles.xml.bz2 and preprocessed it using WikiExtractor. [^6] https://github.com/attardi/wikiextractor We provide the sentence length distribution for the AO-CHILDES and Wikipedia datasets used in this experiment in Appendix B.
@@
We evaluate the grammatical abilities of these models using a developmentally inspired targeted syntactic evaluation benchmark, Zorro Huebner et al. (2021). Zorro is designed for assessing the syntactic and grammatical knowledge of LMs in child-directed language and consists of 13 mid-level categories and 23 subcategories. Each subcategory contains 2,000 sentence pairs, with one grammatically acceptable and one unacceptable sentence per pair. [^7] While Zorro lacks naturalness, this can aid in isolating syntactic ability from lexical or semantic cues Gulordava et al. (2018). Below is an example of a minimal pair from the “Subject-verb agreement (S-v agr)” category: [^8] See Appendix C for the full list of grammatical categories.
@@
By inputting both the acceptable and unacceptable sentences into the model and calculating the proportion of pairs where the model assigns a higher probability to the acceptable sentence, we obtain the grammaticality judgment score (Accuracy). In this study, we report scores for each mid-level category (henceforth, grammatical items) as well as their macro-average.
@@
Table 3: Performance difference when changing the direction of the cognitive constraints in DynamicLimit-Exp. and indicate items where DynamicLimit-Exp (↓) performed significantly better or worse than DynamicLimit-Exp (↑), respectively (z-test for proportions, p < 0.05).
@@ 4.3 Results
Table 2 presents the accuracy of models trained on Wikipedia, showing trends similar to those observed in Table 1, where the models were trained on AO-CHILDES. Specifically, DynamicLimit-Linear and DynamicLimit-Exp outperform NoLimit and StaticLimit in overall accuracy, with DynamicLimit-Exp achieving the highest performance, further supporting the efficacy of incorporating developmental working memory constraints. These findings suggest that the CP effect does not depend solely on exposure to specific linguistic stimuli (e.g., CDS) but rather on the learning algorithm itself, which mirrors human cognitive development. This result aligns with existing research Feng et al. (2024); Padovani et al. (2025), which have reported that CDS is not uniquely valuable for training LMs. This finding suggests that our method is applicable to LLM pretraining, as they typically use non-CDS datasets such as Common Crawl and Wikipedia Touvron et al. (2023a).
Figure 3: Accuracy trajectories over training epochs (1, 5, 10) for six grammatical categories that showed significant final-stage improvements with developmental constraints.
@@ 5.1 Testing the “Less-is-more” Hypothesis with Reversed Cognitive Constraints
[…]
Mark fixed one worn canal, and Roger
  fixed more.
*Mark fixed one canal, and Roger fixed
5.2 Tracking Developmental Gains Across Training
To more directly support our claim that developmentally guided learning simulates a CP, we examine how model performance unfolds over time—not just at the endpoint but at intermediate stages as well. This addresses the need for stage-by-stage comparisons raised by prior evaluations.
Figure 3 tracks accuracy at Epochs 1, 5, and 10 for six grammatical categories selected based on statistically significant improvements observed in Table 1. At the early stage (Epoch 1), models with larger or fixed memory (NoLimit, StaticLimit) perform better. However, the developmentally constrained model, DynamicLimit-Exp, shows steady gains over time, ultimately surpassing these baselines by Epoch 10 in multiple categories. The improvement is especially pronounced in CASE and FILLER.GAP, highlighting a pattern of late-stage acceleration. These results suggest that incrementally increasing memory capacity over training acts as a beneficial inductive bias, enabling the model to generalize more effectively from limited early experience—consistent with the hypothesized role of a CP in human language acquisition.
(a) NoLimit
(b) DynamicLimit-Exp
@@ 5.2 Tracking Developmental Gains Across Training
5.3 Learning to Represent: The Cognitive Effect of Memory Expansion
We analyze representational change by tracking embedding diversity within epochs and shifts between epochs. Specifically, we consider two complementary aspects of representational change: (i) the diversity or dispersion of embeddings within each epoch, which reflects the isotropy and expressiveness of the representation space at a given time; and (ii) the amount of shift in embeddings between epochs, which captures the degree of representational update and learning progress over time.
Figure 4 visualizes t-SNE projected embeddings for the FILLER.GAP category, where DynamicLimit-Exp showed clear gains over baselines (§4.3, §5.1). In the NoLimit model (Figure 4a), embedding clusters expand from Epoch 1 to 5 but subsequently contract and overlap by Epoch 10. This pattern reflects a reduction in within-epoch diversity (i.e., lower isotropy) and minimal between-epoch shift, indicating that the representations stagnate and fail to evolve structurally as training progresses. In contrast, DynamicLimit-Exp (Figure 4b) produces more structured trajectories: clusters remain well-separated within epochs and continue to shift meaningfully across epochs. This suggests not only sustained representational plasticity but also a finer-grained encoding of syntactic distinctions over time.
To quantify these trends, Table 4 reports entropy (capturing embedding dispersion) and mean Euclidean distance between clusters (capturing separation). [^11] The Appendix D shows how to calculate each measure. The NoLimit model shows a drop in entropy and a plateau in inter-cluster distance after Epoch 5, consistent with representational collapse. Meanwhile, DynamicLimit-Exp maintains higher entropy and exhibits a steady increase in distance, consistent with ongoing structural refinement. These findings indicate that developmentally guided memory expansion helps preserve expressive, isotropic, and well-separated embedding spaces—properties that support better generalization in language models Diehl Martinez et al. (2024). [^12] Similar trends were found for CASE; see Appendix E.